# Continuous-time quantum walk-based ansätze on neutral atom hardware

### v1

@@ Abstract
A key use of near-term quantum computers is as analogue simulators, matching the action of some virtual or abstracted system onto a program executed on physical quantum hardware. This work explores continuous-time quantum walks (CTQW) on constrained graphs of independent set configurations and implements these abstract walks on analog-mode neutral-atom hardware. First, we explore variational state preparation protocols with limited controls through the lens of CTQW and optimal control to prepare nontrivial target states in non-separable Hilbert spaces. Next, we match these virtual walk dynamics to physical execution on analog-mode neutral atom hardware by leveraging the Rydberg blockade phenomenon. We analyze the convergence and scaling in the preparation of these known states as an indicator of the upper bounds on the quantum speedup mechanisms predicted by ideal CTQWs, showing that these signatures persist even when executed on noisy hardware. Finally, we introduce noise mitigation methods using Bayesian postprocessing. This paper demonstrates the ability to prepare nontrivial entangled states using quantum walks in constrained subspaces, and that nonequilibrium dynamics on constrained independent set supspaces is feasible on cloud-accessible analog-mode neutral-atom quantum computers.
@@ I Introduction
Quantum search and optimization are key applications for near-term quantum computers [1]. Algorithms for unstructured or spatial search and the Quantum Approximate Optimization Algorithm (QAOA) exemplify the unique properties of quantum computing: superposition to explore exponentially many computational pathways in quantum parallel, entanglement to encode a cost landscape, and interference to amplify desired states [2, 3]. These algorithms typically utilize an alternating sequence of unitaries, where the first diagonal unitary phase-encodes the cost or validity of each solution state, and the second mixing unitary drives amplitude transfer between these states. This alternating structure can be naturally understood through the framework of continuous-time quantum walks (CTQWs), where coherent dynamics of quantum states traversing a graph underpin the convergence behavior of the algorithm.
CTQWs are a foundational concept in quantum algorithm design, providing a universal platform for computation and achieving exponential speedup over classical methods in some instances [4, 5]. Indeed, extension of the original QAOA framework of unconstrained optimization over binary variables to constrained problems and problem spaces that map to higher-order permutation groups has found that CTQW-inspired mixers can offer highly-improved convergence and be less prone to vanishing gradients, by restricting the search to a space of valid solutions or capturing key features of the cost landscape [6, 7, 8, 9].
More generally, CTQW models operate in a non-adiabatic regime and can saturate the Mandelstam-Tamm quantum speed limit [10, 11], making them especially well-suited for describing and analyzing variational heuristics driven by non-equilibrium dynamics [12] and shortcuts to adiabaticity [13].
@@
Experimentally, CTQWs have been realized in the position basis across a variety of platforms, including walks over line graphs using photonic waveguides [14], neutral atoms [15], and trapped ions [16], as well as two-dimensional grid graphs using superconducting qubit arrays [17, 18], and [19] implementing a spatial quantum search on large-scale two-dimensional lattices of strontium atoms. Beyond spatial walks, variational circuits that implement CTQWs in the computational basis have been demonstrated with photonic waveguides [20], and CTQW-like dynamics have been observed in quantum many-body scar dynamics in Rydberg chains [21] on neutral atom systems [22].
@@
In this work, we implement CTQW based ansätze for constrained combinatorial spaces on QuEra’s Aquila neutral-atom platform [23] by mapping abstract walk generators onto the Rydberg Hamiltonian under blockade constraints. We focus on the preparation of two classes of benchmark states–computational-basis product states and coherent states in the symmetric subspace of the walk generator–and analyze their convergence properties under CTQW-informed variational optimization schemes. By comparing ideal CTQW dynamics, noiseless Rydberg emulation, and experimental results, we assess the extent to which characteristic amplification and interference effects of efficient CTQW-based ansätze survive on noisy intermediate-scale hardware. In Section II, we introduce the phase-walk ansätze, performance metrics used in later analysis, the representative target states, and the variational optimization strategies. In Section III, we describe the mapping of the ansätze to the Rydberg Hamiltonian on Aquila. Results of state preparation on hardware are compared with the scaling predicted by ideal CTQW dynamics and noiseless Rydberg emulation in Section IV. In Section V, we apply quenches on the prepared states as an indicator of coherent state preparation. Finally, in Section VI, we summarize our conclusions. Overall, our results establish constrained-subspace CTQWs as a framework for entangled state preparation on analog-mode neutral-atom processors, and demonstrate that CTQW-based ansätze for known states can exhibit the super-quadratic convergence characteristic of efficient quantum-walk protocols.
@@ II Phase-walk ansätze
Together the two unitaries form a parameterized phase-walk ansatz for which a common task is to prepare some target or resource state by alternating between the walk (or “mixer”) generator $\hat{\mathcal{G}}$ and phasor generator $\hat{C}$ applied to some fiducial initial state $|\psi_{0}\rangle$
@@
The phasor generator can be written as some function $c_{a}\equiv f(z)$, typically a low-order polynomial to conform to the locality constraints of quantum hardware. Here, we choose a first-order function $f(z)=\sum_{i}c_{i}z_{i}$ so that the phasor generator can be efficiently written in terms of Pauli matrices
@@
II.1 Walks in constrained subspaces
@@ II.2 State preparation
II.3 Target states
For the constraint graph of the $N$ vertex ring, we chose two classes of target states. The first class are product states, e.g. $|\psi\rangle=|0010101\rangle$ represented by some bitstring $z$. While these states are trivial and unentangled, it may be nontrivial to prepare such a state with a quantum walk, given the constrained global mixer term that can only be broken by a locally tuned value of the phasor $\gamma_{i}$.
The second class consists of superpositions over the symmetric subspace. An $N$-vertex ring has both translational and reflective symmetries invariant under the dihedral group $D_{N}$; accordingly, these symmetry groups coincide with the automorphism group of the constraint graph $G$ – the set of vertex relabelings that preserve the graph’s adjacency. The symmetric subspace is therefore spanned by equal-weight superpositions of all states labelled by bitstrings related by cyclic rotations or reflections,
@@ II.3 Target states
where $z]=\{g\!\cdot\!z:\,g\in D_{N}\}$ is the dihedral orbit of a representative bitstring zz [^1] The first instance in which the symmetric subspace contains states whose entire orbit cannot all be generated by rotations alone occurs at $N=9$, where $\big{|}\big{[}010010001\big{]}\big{\rangle}=\,\tfrac{1}{\sqrt{18}}\big{(}|010010001\rangle+|010010100\rangle+|010001010\rangle+$ $|010001001\rangle+|010100100\rangle+|010100010\rangle+|101001000\rangle+$ $|101000100\rangle+|100101000\rangle+|100100010\rangle+|100010100\rangle+$ $|100010010\rangle+|001010001\rangle+|001010010\rangle+|001001010\rangle+$ $001000101\rangle+|000101001\rangle+|000100101\rangle\big{)}.$ .
We refer to these states as bracelet states, borrowing from the combinatorial term for equivalence classes of bitstrings under cyclic rotations and reflections [30].
II.4 Ansätze performance
@@
II.5 Variational optimization
II.5.1 Product states
@@
Setting the $X$-rotation angles to $(\tau_{0},\tau_{1})\equiv\{\pi/4,\,3\pi/4\}\pmod{2\pi}$ (in any order) and choosing $\gamma=\pm\pi\pmod{2\pi}$ and $c_{k}=z_{k}^{*}$ enables the preparation of any computational-basis state [^2] Geometrically, the single-qubit gates perform an $X\text{-}Z\text{-}X$ rotation sequence on the Bloch sphere of each qubit. Starting from +z+z (i.e., $|0\rangle$), an $X$-rotation by $\tau_{0}=\pi/4$ moves the Bloch vector to y-y. Then, if $c_{k}=1$, the intermediate $Z$-rotation moves the vector to y-y. Finally, an $X$-rotation by $\tau_{1}=3\pi/4$ drives vectors on +y+y back to +z+z (i.e., $|0\rangle$) and those on y-y to z-z (i.e., $|1\rangle$).. However, when a nontrivial connected constraint graph is introduced, each $\sigma_{x}^{k}$ becomes a neighbor-conditioned multi-qubit gate, so the evolution is no longer separable. Nevertheless, the walk ansatz can still prepare the exact state. Given a phasor unitary
@@
as a sum of Pauli-z terms that are non-zero in the target bitstring, three protocols are possible. The first is a Trotterized adiabatic protocol that uses perturbatively small parameters $\gamma,\tau\ll 1$, exact in the $p\to\infty$ limit [35]. The second is a Trotterized mask protocol with $\gamma,\tau\ll 1$ but $\gamma\gg\tau$ to “pin” the zero qubits in the zero state with a large effective Z field. The third is a fast-forward limit [13] with $\gamma=\pi$ and $\tau_{0}+p\tau_{1}=\tfrac{\pi}{2}$, which can be shown to be exact by the following argument. First, the walk Hamiltonian can be split by $U_{z^{*}}$ into two components:
@@
$|\tau_{0},\tau_{1}\rangle=\big{[}R_{-}(\tau_{1})R_{+}(\tau_{1})\big{]}^{p^{\prime}}|\psi_{0}\rangle.$ (20)
By BCH expansion to second-order, a single pair is,
@@
where $S_{j}$ is the set of bitstrings in the positive subspace with Hamming weight $j$ (so $|v_{0}\rangle=|0\rangle$ and $|v_{k}\rangle=|z^{*}\rangle$). The effective Hamiltonian $J$ is the SU(2) spin-$k/2$ chain Hamiltonian (permutation-symmetric noninteracting spins 1/2) expressed in the symmetric Dicke basis, which has non-zero entries only on the first off-diagonals:
@@
since each state in $S_{j}$ has kjk-j forward neighbors in $S_{j+1}$ and each state in $S_{j+1}$ has $j+1$ backward neighbors in $S_{j}$. The lowest-order coupling between $|0\rangle$ and $|z^{*}\rangle$ then occurs at order $k$ with strength,
@@
and, since $J$ inherits the bipartite structure of $\mathcal{G}$,
@@
Figure 2: Success probability of preparing product states $|z^{*}\rangle=|0^{N-2h}(01)^{h}\rangle$, where $h=1$ to $\lfloor N/2\rfloor$, using optimized walk times $\tau_{0}$ and $\tau_{1}$, for systems sizes $N=5$ to $23$. The top-left panel shows success probability across ansatz depths $p=1$ to 55 colored by Hamming weight ($h$), while the bottom-left panel highlights the reference states $z^{\rm half}$ and $z^{\rm MIS}$ (see Section II.3). The right panels show success probability as a function of $\tau_{0}$ and $\tau_{1}$ for the $N=9$ instance of $z^{\rm half}$ at depths $p=1$, 22, and $3$. The optimal points (given in Table 2) are marked with a red ‘X’, and theoretical predictions from Eqs. 31 and 32 is indicated by a yellow ‘+’.
Figure 2 shows the result of running Nelder-Mead from the analytic starting point $(\tau_{0}^{*},\tau_{1}^{*})$ for system sizes $N=5$ to $23$ at depths $p=1$ to $5$. In every case the solver converges to an optima in a single basin. For the largest system $N=23$, the worst-case success probability over all bit-string targets at $p=1$ is only 0.267 (occurring at Hamming weight $h=8$), but this minimum probability rises to 0.925 by $p=5$. Conversely, the $h=11$ state achieves a probability of 0.901 at $p=1$ and essentially unity at $p=5$. At fixed $N$ the probability profile as a function of Hamming weight is well described by a skewed quadratic, with its lowest point near $h\approx\lfloor N/4\rfloor$ and monotonic increase towards both the all-zero and maximum Hamming weight. The $(\tau_{0},\tau_{1})$ for the optima depicted in the right panel are listed in 2. Over $p=1$ to $3$ the effective walk time $T_{\rm{eff}}$, as given by Eq. 28, approaches $\pi/2$.
II.5.2 Bracelet states
The optimization strategy for preparation of bracelet states is guided by two practical considerations, as expanded in Section III. First, we aim to minimize the total evolution time and, second, utilize the hardware-native realization of the phasor unitary

### v2

@@ Abstract
Continuous-time quantum walks offer provable speedups for certain computational problems, yet translating these advantages to near-term hardware remains challenging. We present the first experimental demonstration of variational ansätze based on continuous-time quantum walks on an analog neutral-atom processor. For unentangled targets, we derive closed-form expressions for near-optimal control parameters that transfer directly to hardware with minimal calibration. Experiments on QuEra’s Aquila processor provide the first observation of the super-quadratic convergence characteristic of efficient quantum walk algorithms, visible at low circuit depth, with theory predicting stronger speedups as hardware improves. For entangled targets, specifically symmetric superpositions in the Rydberg-blockaded subspace, we introduce an optimization protocol exploiting spectral properties of the walk dynamics. The required evolution time scales inversely with the spectral gap, offering an advantage over adiabatic protocols that scale to the square of the spectral gap. We demonstrate this scaling behavior on Aquila and verify that the prepared states are coherent superpositions via quench dynamics. This constitutes the first preparation of such symmetric entangled states on neutral-atom hardware. Our results establish a practical pathway from abstract quantum walk algorithms to analog quantum processors, demonstrating that the dynamics underlying their potential for super-quadratic quantum speedup are accessible on current devices.
@@ I Introduction
Quantum search and optimization are key applications for near-term quantum computers [1]. Algorithms for unstructured or spatial search and the Quantum Approximate Optimization Algorithm (QAOA) exemplify the unique properties of quantum computing: superposition to explore exponentially many computational pathways in quantum parallel, entanglement to encode a cost landscape, and interference to amplify desired states [2, 3]. These algorithms typically utilize an ansatz [^1] A parameterised family of circuits used to prepare candidate states, with the parameters chosen (typically by a classical outer loop) to maximize the probability of sampling desirable solutions. comprised of an interleaved sequence of unitaries. At each ansatz iteration, the first diagonal unitary phase-encodes the cost or validity of each solution state, and the second mixing unitary drives amplitude transfer between these states. This layered structure can be naturally understood through the framework of continuous-time quantum walks (CTQWs), where coherent dynamics of quantum states traversing a graph underpin the convergence behavior of the algorithm.
CTQWs provide a natural framework for quantum algorithm design, with coherent dynamics over a graph underlying methods for spatial search and optimization [4, 5]. In particular, CTQW-based search algorithms achieve the optimal quadratic improvement in query complexity for unstructured search [3], and operate in a non-adiabatic regime that can saturate fundamental quantum speed limits [6, 7].
These algorithms can all be expressed as a phase-walk ansatz, an alternating sequence of two interleaved unitaries. With each ansatz iteration a phase separator applies rotations that encode the problem at hand, and a mixing unitary carries out a CTQW that drives amplitude transfer between states. While the query-complexity bound for unstructured search is quadratic, the phase-encoding layer can carry additional information about global problem structure. When symmetries of the mixing graph (whose edges define allowed transitions) reflect the structure of the solution space, compounding interference across symmetric pathways can coherently focus amplitude on target states, yielding super-quadratic convergence [8, 9]. Mixing graphs that enable such “shortcut” dynamics are especially significant for near-term quantum devices, where practical advantage requires sufficient amplification within time-limited coherence windows.
@@
Experimentally, CTQWs have been realized in the position basis across a variety of platforms, including walks over line graphs using photonic waveguides [10], neutral atoms [11] and trapped ions [12], as well as two-dimensional grids using superconducting qubit arrays[13, 14] and tweezer-programmable strontium-atom lattices [15]. Beyond spatial walks, variational circuits that implement CTQWs in the computational basis have been demonstrated with photonic waveguides [16], and CTQW-like dynamics have been observed in quantum many-body scar dynamics in Rydberg chains [17] on neutral atom systems [18]. Full phase-walk ansätze-alternating phase-encoding and mixing unitaries-have been implemented for QAOA on superconducting processors [19] and trapped ions [20].
Despite this progress, each approach addresses only part of the phase-walk framework. Position-basis walk experiments demonstrate coherent quantum dynamics but cannot leverage the full computational-basis Hilbert space and, as the walk graph is fixed by hardware geometry, are not easily tailored to match the structure of a target problem. Gate-based QAOA implementations on superconducting and trapped-ion processors do include phase encoding. However, the mixing unitary is typically limited to a transverse-field operator on an unconstrained Hilbert space, requiring either penalty terms or post-selection to enforce feasibility constraints-both of which incur overhead that grows with system size [21]. Recent QAOA studies have reported evidence of scaling advantages over classical algorithms, but these claims derive entirely from noiseless classical simulations, with hardware experiments demonstrating only Grover-like amplification at low circuit depth [22, 20].
Analog neutral-atom platforms provide an ideal environment for the exploration of phase-walk ansätze. The blockade interaction, whereby strong Rydberg-state repulsion suppresses simultaneous excitation of nearby atoms, enables efficient implementation of constrained CTQWs whose walk graphs possess non-trivial topology with direct relevance to constrained combinatorial optimization [21, 23, 24]. The analog evolution has the potential to natively implement continuous-time walk dynamics, while global and local detuning provide flexible means to apply site-dependent phase shifts encoding the cost function. This combination yields a hardware-efficient realization of the full phase-walk ansatz in which both constraint enforcement and the mixing unitary are “free” in terms of gate overhead.
In this work, we report the first experimental implementation of continuous-time quantum walk (CTQW)-based variational ansätze on analog-mode neutral-atom hardware, using QuEra’s Aquila processor [25]. Unlike prior Rydberg-atom experiments that sample low-energy configurations or observe emergent many-body dynamics [26, 18], we target the preparation of specific, known benchmark states and systematically quantify convergence as a function of ansatz depth. This provides a controlled setting that isolates the intrinsic performance of CTQW-based dynamics from problem-dependent complexity.
We focus on two classes of target states: computational-basis product states, and “bracelet” states–equal-weight superpositions over all basis states related by the dihedral symmetries of the walk graph. The former relies on dynamics that are analogous to the preparation of an unknown unentangled state in search or optimization. The latter generalize Dicke states, the symmetric entangled states widely used in quantum metrology and constrained optimization [27, 21], to an independent-set subspace. While Dicke states span a subspace that grows only linearly with system size, bracelet states live in an exponentially large symmetric sector, making their preparation a more demanding test of coherent control. Futhermore, their preparation probes whether the non-adiabatic dynamics for preparation of product states in constrained Hilbert spaces [28] via quantum many-body scar dynamics [18, 17] extend to the preparation of similarly constrained entangled targets.
For product-state preparation, we derive closed-form expressions for near-optimal walk parameters, providing analytically guided initial points via a method that generalizable to walk graphs defined by independent-set subspaces. For bracelet states, we introduce an optimization protocol based on the spectral properties of the walk dynamics [29]. By comparing ideal CTQW predictions, noiseless Rydberg emulation, and experimental measurements, we identify the depth at which super-quadratic amplification persists and the crossover beyond which noise suppresses convergence, providing a practical figure of merit for NISQ implementations of CTQW-based algorithms. Finally, we employ quench dynamics under the walk Hamiltonian to probe coherence of the prepared bracelet states, providing experimental evidence that the observed populations arise from coherent superposition rather than an incoherent mixture.
@@
The remainder of this work is as follows. In Section II, we introduce the phase-walk ansätze, performance metrics used in later analysis, the representative target states, and the variational optimization strategies. In Section III, we describe the mapping of the ansätze to the Rydberg Hamiltonian on Aquila. Results of state preparation on hardware are compared with the scaling predicted by ideal CTQW dynamics and noiseless Rydberg emulation in Section IV. In Section V, we apply quenches on the prepared states as an indicator of coherent state preparation. Finally, in Section VI, we summarize our conclusions. Overall, our results establish constrained-subspace CTQWs as a framework for entangled state preparation on analog-mode neutral-atom processors, and demonstrate that CTQW-based ansätze for known states can exhibit the super-quadratic convergence characteristic of efficient quantum-walk protocols on currently available hardware.
@@ II.1 Quantum Walk-based Ansẗze
Together the two unitaries form a parameterized phase-walk ansatz for which a common task is to prepare some target or resource state by alternating between the walk (or “mixer”) generator $\hat{\mathcal{G}}$ and phase-separator generator $\hat{C}$ applied to some fiducial initial state $|\psi_{0}\rangle$
@@
The phase-separator generator can be written as some function $c_{a}\equiv f(z)$, typically a low-order polynomial to conform to the locality constraints of quantum hardware. Here, we choose a first-order function $f(z)=\sum_{i}c_{i}z_{i}$ so that the phase-separator generator can be efficiently written in terms of Pauli matrices
@@
II.2 Walks in constrained subspaces
@@ II.3 State preparation
II.4 Target states
For the constraint graph of the $N$-vertex ring, we chose two classes of target states. The first class are product states, e.g. $|\psi\rangle=|0010101\rangle$ specified by a bitstring $z$. While preparing such states for a system of non-interacting qubits is trivial (see Section II.6), preparing these known states using walk-based dynamics provides a benchmark for the typical goal of variational ansätze: preparing unknown, low-entanglement states corresponding to low-cost solutions [2, 35, 21]. Furthermore, because the mixer allows only constraint-respecting transitions, it is not a priori clear that arbitrary product states can be efficiently prepared with high fidelity using an ansatz parameterised solely by walk times and locally tuned phase-separator values $\gamma_{i}$. We focus on two representative product-state targets. The first, denoted $z^{\rm half}$, is the state $|0^{N-2h}(01)^{h}\rangle$ with $h=\lfloor N/2\rfloor$, corresponding to roughly half the maximum Hamming weight for the ring graph. The second, denoted $z^{\rm MIS}$, is a maximum independent set of the ring: the alternating pattern $(01)^{N/2}$ for even $N$, or $0(10)^{(N-1)/2}$ for odd $N$.
The second class consists of superpositions over the symmetric subspace. An $N$-vertex ring has translational and reflective symmetries described by the dihedral group $D_{N}$ (i.e., the set of all discrete rotations and reflections). Because the constraint graph $G$ is defined on the ring, these operations relabel vertices without changing which pairs are adjacent (which are referred to in graph theory as automorphisms of $G$). We therefore consider the corresponding symmetry-invariant subspace, spanned by equal-weight superpositions of all computational-basis states labeled by bitstrings related by cyclic rotations or reflections,
@@ II.4 Target states
where $z]=\{g\!\cdot\!z:\,g\in D_{N}\}$ is the dihedral orbit of a representative bitstring $z$  [^2] The first instance in which the symmetric subspace contains states whose entire orbit cannot all be generated by rotations alone occurs at $N=9$, where $\big|\big[010010001\big]\big\rangle=\,\tfrac{1}{\sqrt{18}}\big(|010010001\rangle+|010010100\rangle+|010001010\rangle+$ $|010001001\rangle+|010100100\rangle+|010100010\rangle+|101001000\rangle+$ $|101000100\rangle+|100101000\rangle+|100100010\rangle+|100010100\rangle+$ $|100010010\rangle+|001010001\rangle+|001010010\rangle+|001001010\rangle+$ $001000101\rangle+|000101001\rangle+|000100101\rangle\big).$ .
We refer to these states as bracelet states, borrowing from the combinatorial term for equivalence classes of bitstrings under cyclic rotations and reflections [36]. The preparation of such states is an informative benchmark, as it requires coherent many-body interference to establish uniform amplitudes across a dihedral orbit of up to $2N$ basis states. Futhermore, bracelet states have potential applications as initial states for constrained optimization [21, 37] and quantum metrology [27].
II.5 Ansätze performance
@@
II.6 Variational optimization
II.6.1 Product states
@@
Setting the $X$-rotation angles to $(\tau_{0},\tau_{1})\equiv\{\pi/4,\,3\pi/4\}\pmod{2\pi}$ (in any order) and choosing $\gamma=\pm\pi\pmod{2\pi}$ and $c_{k}=z_{k}^{*}$ enables the preparation of any computational-basis state [^3] Geometrically, the single-qubit gates perform an $X\text{-}Z\text{-}X$ rotation sequence on the Bloch sphere of each qubit. Starting from +z+z (i.e., $|0\rangle$), an $X$-rotation by $\tau_{0}=\pi/4$ moves the Bloch vector to y-y. Then, if $c_{k}=1$, the intermediate $Z$-rotation moves the vector to y-y. Finally, an $X$-rotation by $\tau_{1}=3\pi/4$ drives vectors on +y+y back to +z+z (i.e., $|0\rangle$) and those on y-y to z-z (i.e., $|1\rangle$).. However, when a nontrivial connected constraint graph is introduced, each $\sigma_{x}^{k}$ becomes a neighbor-conditioned multi-qubit gate, so the evolution is no longer separable. Nevertheless, the walk ansatz can still prepare the exact state. Given a phase separator
@@
as a sum of Pauli-z terms that are non-zero in the target bitstring, three protocols are possible. The first is a Trotterized adiabatic protocol that uses perturbatively small parameters $\gamma,\tau\ll 1$, exact in the $p\to\infty$ limit [42]. The second is a Trotterized mask protocol with $\gamma,\tau\ll 1$ but $\gamma\gg\tau$ to “pin” the zero qubits in the zero state with a large effective Z field. The third is a fast-forward limit [43] with $\gamma=\pi$ and $\tau_{0}+p\tau_{1}=\tfrac{\pi}{2}$, for which we give the following argument. First, the walk Hamiltonian can be split by $U_{z^{*}}$ into two components:
@@
$|\tau_{0},\tau_{1}\rangle=\big[R_{-}(\tau_{1})R_{+}(\tau_{1})\big]^{p^{\prime}}|\psi_{0}\rangle,$ (20)
where $|\psi_{0}\rangle$ is defined as in Eq. 5. By BCH expansion to second-order, a single pair is,
@@
where $S_{j}$ is the set of bitstrings in the positive subspace with Hamming weight $j$, and the $|v_{j}\rangle$ are the corresponding uniform “shell” states of fixed Hamming weight (so $|v_{0}\rangle=|0\rangle$ and $|v_{k}\rangle=|z^{*}\rangle$). The effective Hamiltonian $J$ is the SU(2) spin-$k/2$ chain Hamiltonian (permutation-symmetric noninteracting spins $1/2$) expressed in the symmetric Dicke basis, so that in this subspace it reduces to a one-dimensional ladder $\{|v_{j}\rangle\}_{j=0}^{k}$, which has non-zero entries only on the first off-diagonals:
@@
since each state in $S_{j}$ has kjk-j forward neighbors in $S_{j+1}$ and each state in $S_{j+1}$ has $j+1$ backward neighbors in $S_{j}$. Because $J$ only couples neighboring Hamming-weight shells $j\leftrightarrow j+1$, the first non-vanishing contribution connecting $|v_{0}\rangle$ to $|v_{k}\rangle$ appears at order $k$. The lowest-order coupling between $|0\rangle$ and $|z^{*}\rangle$ then occurs at order $k$ with strength,
@@
Moreover, since $J$ inherits the bipartite structure of $\mathcal{G}$, the transition amplitude $\langle v_{k}|e^{-iT_{\rm eff}J}|v_{0}\rangle$ has a fixed parity in $T_{\rm eff}$ determined by $k$, and, since $J$ inherits the bipartite structure of $\mathcal{G}$,
@@
Figure 2: Success probability of preparing product states $|z^{*}\rangle=|0^{N-2h}(01)^{h}\rangle$, where $h=1$ to $\lfloor N/2\rfloor$, using optimized walk times $\tau_{0}$ and $\tau_{1}$, for systems sizes $N=5$ to $23$. Top-left shows success probability across ansatz depths $p=1$ to 55 colored by Hamming weight ($h$). Bottom-left highlights the reference states $z^{\rm half}$ and $z^{\rm MIS}$ (see Section II.4). Right shows success probability as a function of $\tau_{0}$ and $\tau_{1}$ for the $N=9$ instance of $z^{\rm half}$ at depths $p=1$, 22, and $3$. The optimal points (given in Table 2) are marked with a red ’X’, and theoretical predictions from Eqs. 31 and 32 is indicated by a yellow ’+’.
Figure 2 shows the result of running Nelder-Mead from the analytic starting point $(\tau_{0}^{*},\tau_{1}^{*})$ for system sizes $N=5$ to $23$ at depths $p=1$ to $5$. In every case the solver converges to an optima in a single basin. For the largest system $N=23$, the worst-case success probability over all bit-string targets at $p=1$ is only 0.267 (occurring at Hamming weight $h=8$), but this minimum probability rises to 0.925 by $p=5$. Conversely, the $h=11$ state achieves a probability of 0.901 at $p=1$ and essentially unity at $p=5$. At fixed $N$ the probability profile as a function of Hamming weight is well described by a skewed quadratic, with its lowest point near $h\approx\lfloor N/4\rfloor$ and monotonic increase towards both the all-zero and maximum Hamming weight. The $(\tau_{0},\tau_{1})$ for the optima depicted in the right panel are listed in 2. Each of the three depicted instances show $(\tau_{0}^{*},\tau_{1}^{*})$ within the locally convex region of the optima. Over $p=1$ to $3$ the effective walk time $T_{\rm{eff}}$, as given by Eq. 28, approaches $\pi/2$.
II.6.2 Bracelet states
The optimization strategy for preparation of bracelet states is guided by two practical considerations, as expanded in Section III. First, we aim to minimize the total evolution time and, second, utilize the hardware-native realization of the phase separator